OpenStudy (anonymous):

Will medal An ice cream store is making waffle cones. The finished cone should have a radius of 1 inch and be 5 inches tall. The slant height of the cone is 5.1 inches. How much waffle will they need to make each cone A. 6 in2 B. 16 in2 C. 18 in2 D. 19 in2

OpenStudy (trojanpoem):

 find the area of the cone.

OpenStudy (trojanpoem):

Area = pi * r * slant = 3.14 * 1 * 5.1 = 16 in^2 (don't add pir^2 ) as it doesn't have a cover.

OpenStudy (anonymous):

thanks i got it now :)

OpenStudy (trojanpoem):

any time.